# Playing Language Game with LLMs Leads to Jailbreaking

### v1

@@ 1 Introduction
To validate the effectiveness of our findings, we select four different natural language games and design eight types of custom language games. These two jailbreak attack approaches are tested on GPT-4o, GPT-4o-mini and Claude-3.5-Sonnet using randomly selected safety questions from the SALAD-Bench benchmark (Li et al., 2024). Experimental results show that both natural language games and custom language games can successfully bypass the safety alignment of LLMs, achieving attack success rates of 93% on GPT-4o, 89% on GPT-4o-mini and 83% on Claude-3.5-Sonnet. We also conducted experiments on fine-tuned Llama-3.1-70B to explore the generalization ability of LLMs and discovered that safety alignments achieved through supervised fine-tuning fail to generalize effectively to other domains. These findings suggest that current safety alignment methods do not effectively generalize safety knowledge, leaving substantial room for improvement. Our key contributions can be summarized as follows:
@@ 2 Related Work
Jailbreak Attacks. The safety of LLMs has been a longstanding concern, as adversaries continually develop new methods to manipulate these models into generating harmful content. One key approach to jailbreak attacks, as described by (Wei et al., 2024), involves designing competing objectives within the input. This tactic exploits the LLM’s instruction-following capabilities, pushing the model to prioritize generating a response based on the user’s request rather than adhering to safety constraints. Studies by (Li et al., 2023), (Chang et al., 2024), (Jiang et al., 2024), and (Guo et al., 2024) demonstrate how carefully crafted natural language instructions can trick LLMs into producing harmful or unethical content. By embedding harmful intent in seemingly benign instructions, attackers can confuse the model’s objective and circumvent its safety alignment protocols.
@@
Another prominent jailbreak method involves mismatched generalization, wherein input prompts are encoded in a way that was not accounted for in the LLM’s safety training but still lies within the model’s general pretraining corpus. As noted by (Wei et al., 2024), this technique has been used to bypass alignment processes in various attack schemes. While it provides a comprehensive overview of 28 existing jailbreak strategies, new methods continue to emerge, indicating that the problem is far from solved. For instance, (Deng et al., 2023) demonstrated that interacting with LLMs in medium and low-resource languages can lead to the generation of unsafe outputs. This highlights a significant gap in the models’ safety mechanisms, particularly in under-represented languages that may not have been as rigorously aligned for safety during training. Similarly, (Yuan et al., 2023) showed that conversations encoded with ciphers can trick LLMs into producing unsafe responses. These examples reveal how attackers can use obfuscation strategies to bypass content moderation mechanisms by encoding harmful queries in formats that the models do not immediately recognize as dangerous. In addition, (Liu et al., 2023) proposed AutoDAN, an automated system that generates stealthy jailbreak prompts using a hierarchical genetic algorithm. AutoDAN systematically creates prompts designed to evade safety protocols by mimicking benign inputs, leveraging the evolutionary process to develop increasingly effective jailbreak training strategies over time.
@@
Safety Training of LLMs. Ensuring the responsible and effective deployment of LLMs requires aligning their outputs with human preferences and ethical standards (Korbak et al., 2023; Achiam et al., 2023; Touvron et al., 2023). Common alignment methods include supervised fine-tuning (Bianchi et al., 2023), red-teaming (Ganguli et al., 2022; Perez et al., 2022), the use of reward signals (Ouyang et al., 2022), and preference-based modeling (Christiano et al., 2017). With the rise of reinforcement learning from human feedback (RLHF), (Dai et al., 2023) proposed Safe RLHF, a novel framework which incorporates a two-dimensional human annotation scheme and a safety-focused training mechanism to enhance model performance while ensuring safety. Additionally, (Ji et al., 2024) developed the PKU-SafeRLHF dataset to provide training data and a reproducible code pipeline, facilitating further research in alignment. However, while these methods evaluate the effectiveness of alignment, they do not thoroughly explore whether safety knowledge is generalized across the intermediate layers of LLMs.
@@ 3.1 Natural Language Game Attack
Table 1: The names and the basic rules (Wikipedia contributors, 2024a, b) of the selected four natural language games that operate within the English language.
@@ 4.5 Exploration on Safety Alignment Generalization
To explore the safety alignment generalization abilities of LLMs, we fine-tuned Llama-3.1-70B using the corpus transformed by the custom language games and then conducted custom language game jailbreak attacks on the fine-tuned model. The goal of this process was to evaluate how well the model could generalize its safety alignments to other custom language game variations after fine-tuning on a specific transformation. Specifically, we collected a general knowledge dataset from (Chen et al., 2023) and mixed it with our custom jailbreak dataset. The ratio of the general knowledge dataset to the jailbreak dataset was set at 2.7:1, ensuring a balanced training set that emphasizes both benign content and adversarial examples. For our jailbreak dataset, the output labels were specifically set as negative. To implement the fine-tuning, we utilized the LoRA method (Hu et al., 2021), a parameter-efficient approach for adapting large models by adding low-rank adapters, which reduces computational overhead while preserving the performance of the base model. This method allowed us to effectively fine-tune the Llama-3.1-70B model without retraining the entire network, making the process faster and more resource-efficient.

### v2

@@ 1 Introduction
To validate the effectiveness of our findings, we select four different natural language games and design eight types of custom language games. These two jailbreak attack approaches are tested on GPT-4o, GPT-4o-mini and Claude-3.5-Sonnet using randomly selected safety questions from the SALAD-Bench benchmark Li et al. (2024). Experimental results show that both natural language games and custom language games can successfully bypass the safety alignment of LLMs, achieving attack success rates of 93% on GPT-4o, 89% on GPT-4o-mini and 83% on Claude-3.5-Sonnet. We also conducted experiments on fine-tuned Llama-3.1-70B to explore the generalization ability of LLMs and discovered that safety alignments achieved through supervised fine-tuning fail to generalize effectively to other domains. These findings suggest that current safety alignment methods do not effectively generalize safety knowledge, leaving substantial room for improvement. Our key contributions can be summarized as follows:
@@ 2 Related Work
Jailbreak Attacks. The safety of LLMs has been a longstanding concern, as adversaries continually develop new methods to manipulate these models into generating harmful content. One key approach to jailbreak attacks, as described by Wei et al. (2024), involves designing competing objectives within the input. This tactic exploits the LLM’s instruction-following capabilities, pushing the model to prioritize generating a response based on the user’s request rather than adhering to safety constraints. Studies by Li et al. (2023), Chang et al. (2024), Jiang et al. (2024), and Guo et al. (2024) demonstrate how carefully crafted natural language instructions can trick LLMs into producing harmful or unethical content. By embedding harmful intent in seemingly benign instructions, attackers can confuse the model’s objective and circumvent its safety alignment protocols.
@@
Another prominent jailbreak method involves mismatched generalization, wherein input prompts are encoded in a way that was not accounted for in the LLM’s safety training but still lies within the model’s general pretraining corpus. As noted by Wei et al. (2024), this technique has been used to bypass alignment processes in various attack schemes. While it provides a comprehensive overview of 28 existing jailbreak strategies, new methods continue to emerge, indicating that the problem is far from solved. For instance, Deng et al. (2023) demonstrated that interacting with LLMs in medium and low-resource languages can lead to the generation of unsafe outputs. This highlights a significant gap in the models’ safety mechanisms, particularly in under-represented languages that may not have been as rigorously aligned for safety during training. Similarly, Yuan et al. (2023) showed that conversations encoded with ciphers can trick LLMs into producing unsafe responses. These examples reveal how attackers can use obfuscation strategies to bypass content moderation mechanisms by encoding harmful queries in formats that the models do not immediately recognize as dangerous. In addition, Liu et al. (2023) proposed AutoDAN, an automated system that generates stealthy jailbreak prompts using a hierarchical genetic algorithm. AutoDAN systematically creates prompts designed to evade safety protocols by mimicking benign inputs, leveraging the evolutionary process to develop increasingly effective jailbreak training strategies over time.
@@
Safety Training of LLMs. Ensuring the responsible and effective deployment of LLMs requires aligning their outputs with human preferences and ethical standards (Korbak et al., 2023; Achiam et al., 2023; Touvron et al., 2023). Common alignment methods include supervised fine-tuning (Bianchi et al., 2023), red-teaming (Ganguli et al., 2022; Perez et al., 2022), the use of reward signals (Ouyang et al., 2022), and preference-based modeling (Christiano et al., 2017). With the rise of reinforcement learning from human feedback (RLHF), Dai et al. (2023) proposed Safe RLHF, a novel framework which incorporates a two-dimensional human annotation scheme and a safety-focused training mechanism to enhance model performance while ensuring safety. Additionally, Ji et al. (2024) developed the PKU-SafeRLHF dataset to provide training data and a reproducible code pipeline, facilitating further research in alignment. However, while these methods evaluate the effectiveness of alignment, they do not thoroughly explore whether safety knowledge is generalized across the intermediate layers of LLMs.
@@ 3.1 Natural Language Game Attack
Table 1: The names and the basic rules (Wikipedia contributors, 2024a; b) of the selected four natural language games that operate within the English language.
@@ 4.5 Exploration on Safety Alignment Generalization
To explore the safety alignment generalization abilities of LLMs, we fine-tuned Llama-3.1-70B using the corpus transformed by the custom language games and then conducted custom language game jailbreak attacks on the fine-tuned model. The goal of this process was to evaluate how well the model could generalize its safety alignments to other custom language game variations after fine-tuning on a specific transformation. Specifically, we collected a general knowledge dataset from Chen et al. (2023) and mixed it with our custom jailbreak dataset. The ratio of the general knowledge dataset to the jailbreak dataset was set at 2.7:1, ensuring a balanced training set that emphasizes both benign content and adversarial examples. For our jailbreak dataset, the output labels were specifically set as negative. To implement the fine-tuning, we utilized the LoRA method (Hu et al., 2021), a parameter-efficient approach for adapting large models by adding low-rank adapters, which reduces computational overhead while preserving the performance of the base model. This method allowed us to effectively fine-tune the Llama-3.1-70B model without retraining the entire network, making the process faster and more resource-efficient.